# Token-Budget-Aware LLM Reasoning

### v3

@@ 6.1 Experiment Setup
6.3 RQ2. Effectiveness of Budget Estimation.
6.4 RQ3. Generalization of TALE.
@@
Table 3 demonstrates the generalization of TALE across Yi-lightning Wake et al. (2024), GPT-4o-mini OpenAI (2024a), and GPT-4o OpenAI (2024b) on MathBench-College, showing its ability to reduce output tokens and expenses while maintaining competitive accuracy significantly. TALE achieves substantial token savings, reducing output tokens by 62.6% for Yi-lightning, 61.5% for GPT-4o-mini, and 69.8% for GPT-4o, compared to Vanilla CoT. Expense reductions are equally notable, with costs decreasing from 21.55 to 17.25 for Yi-lightning, 81.56 to 45.60 for GPT-4o-mini, and 1359.42 to 759.95 for GPT-4o. Despite these cost savings, TALE maintains strong accuracy, achieving 76.67% on Yi-lightning, 70.00% on GPT-4o-mini, and 80.00% on GPT-4o, comparable to Vanilla CoT. These results highlight TALE’s effectiveness in balancing cost efficiency and reasoning performance across diverse LLM architectures. The observed accuracy drop is most significant for GPT-4o-mini. This could be attributed to its smaller number of parameters, which makes it more challenging to answer correctly within a limited response reasoning length.
@@ 7 Conclusion
In this paper, we introduce TALE, a framework that reduces token redundancy in Chain-of-Thought (CoT) reasoning by incorporating token budget awareness. TALE dynamically estimates token budgets based on task complexity to guide reasoning, balancing efficiency and accuracy. Experiments show that TALE reduces token usage by 68.9% on average with less than a 5% accuracy loss, outperforming Vanilla CoT in cost-effectiveness while generalizing well across various LLMs.
@@ Appendix A Appendix
Figure 7: An intuitive example to illustrate the workflow of TALE.

### v4

@@ 5.3 TALE Post-Training (TALE-PT)
o3-mini 96.00% 601.51 336.69 97.33% 1163.55 638.46 96.66% 677.65 385.12
Table 4: Comparison of TALE-PT (post-training to internalize token-budget awareness) and other prompt engineering methods. Two different post-training methods, SFT and DPO, are taken into consideration.
Directly Answering Vanilla CoT TALE-PT-SFT TALE-PT-DPO
LLM Dataset ACC $\uparrow$ Output Tokens $\downarrow$ ACC $\uparrow$ Output Tokens $\downarrow$ ACC $\uparrow$ Output Tokens $\downarrow$ ACC $\uparrow$ Output Tokens $\downarrow$
GSM8K 21.00% 38.54 77.56% 241.51 78.57% 139.63 74.11% 149.93
Llama-3.1-8B-Instruct GSM8K-Zero 70.32% 13.49 65.04% 251.08 78.43% 77.85 78.41% 113.41
DPO-based Internalization. Another way to incentivize $\mathcal{M}_{\theta}$ to learn the token-budget preference is applying the DPO algorithm Rafailov et al. (2023) to post-train the model. DPO directly refines the policy through a classification objective, aligning the model’s behavior with the desired preferences. The goal of DPO here is to refine $\mathcal{M}_{\theta}$ so it can accurately solve a given problem ${\bm{x}}$ while adhering to an internalized token budget. We use the target outputs $y_{i}$ from the searched optimal budget as positive samples, while outputs $y_{i}^{\prime}$ generated with the vanilla CoT prompt serve as negative samples. These positive-negative pairs are then used to create the pairwise preference data for DPO training. Given the crafted dataset $\mathcal{D}$ = $\{({\bm{x}}_{i},y_{i},y_{i}^{\prime})\}_{i=1}^{N}$, the objective is to maximize the likelihood that the model ranks the positive samples higher than the negative ones. Formally, we aim to optimize the following objective:
$\mathcal{L}_{\mathrm{DPO}}(\theta)$ $=-\frac{1}{N}\sum_{i=1}^{N}\log P_{\theta}(y_{i}\succ y_{i}^{\prime}),\quad where$
$P_{\theta}(y_{i}\succ y_{i}^{\prime})$ $=\frac{\exp\bigl{(}s(y_{i},\mathbf{x}_{i})\bigr{)}}{\exp\bigl{(}s(y_{i},\mathbf{x}_{i})\bigr{)}+\exp\bigl{(}s(y_{i}^{\prime},\mathbf{x}_{i})\bigr{)}}.$
$P_{\theta}(y_{i}\succ y_{i}^{\prime})$ is the preference function. Here, $s(y_{i},{\bm{x}}_{i})$ is defined as $\sum_{t=1}^{T_{i}}\log\mathbb{P}(y_{i,t}|y_{i,<t},{\bm{x}}_{i})$, and it represents the log-probability of the model generating $y_{i}$ for input ${\bm{x}}_{i}$, which serves as the preference score assigned to $y_{i}$. This score measures how strongly the model favors that output. The objective ensures that the model prioritizes concise and token-efficient outputs while maintaining high-quality reasoning and correctness. During training, the LLM is encouraged to internalize the token budget constraint and adopt a more compact reasoning process guided by the target outputs generated in the first stage. This two-stage process effectively trains the LLM to produce concise yet accurate responses, striking a balance between reasoning quality and token efficiency during inference. More details are in Section A.3.
6 Evaluation
In this section, we provide the experiment results to evaluate the effectiveness of two versions of TALE, TALE-EP and TALE-PT.
6.1 Experiment Setup
Datasets. To evaluate the LLM performance, three most challenging mathematical datasets are taken into consideration: GSM8K Cobbe et al. (2021), GSM8K-Zero Chiang and Lee (2024), and MathBench Liu et al. (2024). GSM8K-Zero, derived from the GSM8K dataset, specifically targets the analysis of over-reasoning and redundancy in LLM-generated outputs. In short, GSM8K-Zero is designed so that the answers are embedded within the questions themselves. LLMs can easily generate correct responses without complicated additional reasoning or redundant calculations.
Models. We conduct experiments on five state-of-the-art LLMs (i.e., GPT-4o OpenAI (2024b), GPT-4o-mini OpenAI (2024a), Yi-lightning Wake et al. (2024), o3-mini OpenAI (2025)), and Lllama-3.1-8B-Instruct Dubey et al. (2024).
Metrics. The target of TALE is to balance the LLM correctness performance and extra redundant token costs. Specifically, TALE seeks to minimize Number of Output Tokens while maintaining comparable Accuracy (Acc) simultaneously.
Number of Output Tokens. We evaluate the token costs by calculating the average output token consumption for each specific task. The output token costs are measured as follows: Number of Output $\textit{Number of Output Tokens}=\frac{1}{N}\sum_{i=1}^{N}{\mathbb{T}(\mathcal{M}({\bm{x}}_{i}))}$, where ${\bm{x}}_{i}$ represents the given question, and $\mathbb{T}$ is a function that measures the number of tokens. Intuitively, the more output tokens, the higher the costs incurred by $\mathcal{M}$. To evaluate costs more precisely, we calculate the average expense per sample. The total token expense includes both input and output tokens used during the query process.
6.2 Effectiveness of TALE-EP
Table 2 compares TALE-EP with other prompt engineering methods across seven datasets, evaluating accuracy, output tokens, and expenses. Effective prompts should maximize accuracy while minimizing token usage and cost. Direct Answering is the most cost-efficient (14.57 tokens, 25.37 expense) but with low accuracy (52.31%). Vanilla CoT achieves the highest accuracy (83.75%) but at a high cost (461.25 tokens, 289.78 expense). TALE-EP balances performance and efficiency, achieving 81.03% accuracy while reducing token usage to 32% and expenses to 41% of Vanilla CoT. On GSM8K, it even surpasses Vanilla CoT with 84.46% accuracy. Note that expense is not directly proportional to output tokens because it also accounts for input and cached tokens. TALE-EP reduces token costs by 68.64% on average, offering a scalable, cost-effective solution for budget-constrained reasoning tasks.
@@ 6.2 Effectiveness of TALE-EP
To further evaluate the generalization of TALE-EP across different LLMs. We conduct experiments across Yi-lightning, GPT-4o-mini, GPT-4o and o3-mini on MathBench-College. Table 3 illustrates the results, showing TALE-EP’s ability to reduce output tokens and expenses while maintaining competitive accuracy significantly. TALE-EP achieves substantial token savings, reducing output tokens by 64.63% on average, compared to Vanilla CoT. Expense reductions are equally notable, with costs decreasing by 45.30% on average. Despite these cost savings, TALE-EP maintains strong accuracy, achieving 76.67% on Yi-lightning, 70.00% on GPT-4o-mini, and 80.00% on GPT-4o, comparable to Vanilla CoT. These results highlight TALE-EP’s effectiveness in balancing cost efficiency and reasoning performance across diverse LLM architectures. The observed accuracy drop is most significant for GPT-4o-mini. This could be attributed to its smaller number of parameters, which makes it more challenging to answer correctly within a limited response reasoning length.
6.3 Effectiveness of TALE-PT
Table 4 compares TALE-PT methods with Vanilla CoT and Direct Answering on GSM8K and GSM8K-Zero using Llama-3.1-8B-Instruct. For GSM8K, Direct Answering demonstrates the lowest token usage (38.54) but at the cost of significantly reduced accuracy (21.00%). In contrast, Vanilla CoT achieves much higher accuracy (77.56%) but incurs a significant increase in token cost (241.51). Note that on GSM8K-Zero, the accuracy of Vanilla CoT drops below Direct Answering. This drop can be attributed to overthinking, as GSM8K-Zero is simpler, with answers often implied directly within the question. In such cases, a long reasoning process can introduce unnecessary complexity, leading to reduced accuracy. Among the TALE-PT methods, TALE-PT-SFT achieves the best accuracy (78.57%, 78.43%) with reduced tokens, while TALE-PT-DPO balances accuracy (74.11%, 78.41%) and token efficiency, cutting token consumption by over 50% on GSM8K-Zero compared to Vanilla CoT.
@@ 7 Conclusion
In this paper, we introduce TALE, a framework that reduces token redundancy in Chain-of-Thought (CoT) reasoning by incorporating token budget awareness. TALE dynamically adjusts the number of reasoning tokens based on the reasoning complexity of each problem, balancing token efficiency and answer correctness. Experiments show that TALE reduces output token usage and expense significantly with acceptable accuracy loss, outperforming Vanilla CoT in cost-effectiveness while generalizing well across various LLMs.
8 Limitations
The experiments of our proposed token-budget-aware reasoning framework currently focus on LLMs that process only text as input and output. While the results demonstrate significant improvements in efficiency and cost reduction, it does not account for models that have multimodal output content. Such as the models generate interleaved images and text as output. In future work, we will extend token-budget awareness to such LLMs with multimodal output by introducing modality-specific budget constraints and designing adaptive strategies to optimize token efficiency for different modality types, such as images and videos.
@@ Appendix A Appendix
Figure 7: An intuitive example to illustrate the workflow of TALE-EP on GPT-4o-mini OpenAI (2024a).
@@ A.1 Definition of Ideal Budget Range
A.2 Effectiveness of Budget Estimation.
A.3 Details of TALE’s Implementation
In this section, we introduce the hyper-parameters used for TALE-EP and TALE-PT.
TALE-EP. TALE-EP uses a zero-shot mechanism to estimate the token budget and then prompts the LLM. The instruction prompts used during this process are shown in Figure 7. To ensure output consistency, we set the temperature to 0.1 and limit the model to a single reasoning path. Additionally, the random seed is fixed at 1024.
TALE-PT. TALE-PT includes two implementations: SFT and DPO. For parameter efficiency, both implementations adopt LoRA Hu et al. (2021) for post-training, with rank set to 8 and lora alpha set to 32. For TALE-PT-SFT, we train for 3 epochs with a batch size of 16, a learning rate of 1e-4, and a weight decay of 0.01. For TALE-PT-DPO, we train for 2 epochs with a batch size of 16, a learning rate of 3e-5, and a weight decay of 0.001.